OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y =x2 - 10x + 25 A. 0 B. 25 C. -25 D. 50 how would you solve this?

OpenStudy (raden):

the quadratic function above can be rewrite as : y = x^2 - 10x + 25 (it is minimum because the coef of x^2 > 0) y = (x - 5)^2 + 0 so, the minimum value is 0 when x = 5

OpenStudy (anonymous):

and the maximum?

OpenStudy (raden):

there two kinds of quadratic fucntion, max and minimum. max when the coefficient of x^2 < 0 and minimum when the coefficient of x^2 > 0 look at the quadratic above, y =x2 - 10x + 25 known the coefficient of x^2 = 1 ( it is minimum, because >0) so, cant be maximum!

OpenStudy (raden):

* there are

OpenStudy (anonymous):

so it would be A right?

OpenStudy (raden):

yes, you are right

OpenStudy (anonymous):

ok thank you very much :)

OpenStudy (raden):

wlcm :)